# TinySQL: A Progressive Text-to-SQL Dataset for Mechanistic Interpretability Research

### v3

@@ Abstract
Abir Harrasse*1  Philip Quirke*1thanks: * Equal contributions. Correspondence to: philip@withmartian.com.  Clement Neo*2 Dhruv Nathawani3Luke Marks1Amir Abdullah4,1 1Martian  2Apart Research  3Gretel.ai  4Cynch.ai
@@ 1 Introduction
Existing text-to-SQL datasets like Spider (Yu et al., 2019) and WikiSQL (Zhong et al., 2017) are too complex and noisy for rigorous interpretability analysis. To address this, we introduce TinySQL, a curated dataset that enables controlled analysis of how transformers learn and generate SQL queries. TinySQL progresses through text-to-SQL tasks of increasing complexity, isolating key aspects of the generation process while maintaining structural consistency. Alongside the dataset, we train and release models of 33M, 0.5B, and 1B parameters, providing a comprehensive testbed for interpretability research.
@@ 2.1 Mechanistic Interpretability Approaches
Recent work with SAEs (Huben et al., 2023; Templeton et al., 2024) has enabled large-scale model behavior analysis, with Marks et al. (2024) using SAE features to uncover interpretable circuits for tasks like subject-verb agreement. However, a key challenge remains: once we understand a model’s behavior in one context, the next meaningful step is unclear, hindering systematic research and clear progress in model mechanism understanding.
@@ 2.3 Text-to-SQL Datasets
Spider (Yu et al., 2019) established a foundation for text-to-SQL research by focusing on cross-domain generalization through multiple tables and schemas. Its variants explore different aspects of complexity: Spider-Syn (Gan et al., 2021) tests semantic understanding by replacing schema terms with synonyms, while Spider-Real (Deng et al., 2020) aligns more closely with natural user queries by omitting explicit column names. These datasets collectively enable research into model robustness across synthetic and real-world scenarios.
The field evolved from simpler beginnings with WikiSQL (Zhong et al., 2017), which focused on single-table queries from Wikipedia to enable large-scale evaluation. More recent efforts like Gretel synthetic SQL (Meyer et al., 2024) and SQL-create-context (b mc2, 2023) have balanced synthetic data generation with authentic query patterns, reflecting the community’s growing emphasis on combining controlled generation with real-world applicability.
However, these existing datasets prioritize training high-performance models through diverse, complex queries rather than supporting controlled experiments. Even synthetic resources like gretelai/synthetic_text_to_sql lack the systematic progression of complexity needed for mechanistic interpretability research. This gap motivates our development of TinySQL, which draws inspiration from machine learning’s long history of using synthetic data to control task complexity and enable clear evaluation. By providing carefully controlled progression of query complexity while maintaining consistent structure, TinySQL bridges the divide between performance-focused datasets and the controlled environment needed for interpretability research.
@@ 3.1 Dataset Structure
The complete dataset consists of nine task variants (CS1/2/3, CS1/2/3/4/5-Syn, and CS1-Nat), each containing 100,000 examples generated through a systematic data creation pipeline that ensures consistency with our design principles.
We also include in Table 9 (see Appendix 7) several statistics on our text questions noted in the work of Mitsopoulou and Koutrika (2025) that analyzes popular Text-To-SQL benchmarks.
Namely, (a) Rarity: the ratio of the rare words to content words of an NL question Read and Read (2000); (b) Lexical density: the ratio of the content words to the total words of an NL question Read and Read (2000), and (c) Flesch reading ease, the ease of reading a sentence Flesch and Gould (1949).
Our datasets have higher lexical density and rarity than most Text-To-SQL datasets including Spiderb mc2 (2023) and BirdLi et al. (2023) (see Figure 8 in Mitsopoulou and Koutrika (2025)), and lower reading ease than all of the Text-To-SQL datasets studied in the same (see Figure 10 of the same work). This is perhaps due to the synthetic nature of our datasets which have a higher concentration of table and field names, and demonstrates that our datasets present a unique challenge with respect to other publicly available resources.
For details on our dataset generation methodology, including schema design, query construction, and instruction templating for CSx, CSx_Syn, and CSx_Nat, please refer to Appendices C.2 and C.3.
@@ 4.4 Sparse Circuit Identification: Method and Results
We train SAEs on the collected activations using the Belrose (2024)’s package, following the top-$k$ training regime Gao et al. (2024) , with $k=16$, an expansion factor of 4, and a learning rate of $8\times 10^{-4}$ for 6 training epochs. Across all models and datasets, we find that the Fraction of Unexplained Variance (FVU) remains below 0.05, indicating that the SAEs faithfully reconstruct the input representations. Training each SAE takes approximately 1 hour for BM1 and 2 hours for BM2, all on a single A100 GPU.
@@
To isolate task-relevant features, we follow the four-step procedure in Algorithm 1. We evaluate each circuit’s fidelity using zero-ablation and report recovered accuracy in Table 2, with additional results on other combinations in Appendix I.
@@
Table 2: Recovered accuracy and layer-wise head selections for circuits identified via high-AUC SAE features. Each layer column is split into two rows per model (first row for attention heads and second for MLP neurons). Full results are in Appendix H.
@@ 5.1 Identifiability via SAE-EAP Overlap
Model Dataset Layer 1 Layer 2
@@ 5.2 Local Minimality of Circuits
Model Dataset Percentage of Model Used
@@
We also explored an alternative circuit extraction strategy based on mean ablation . While this approach produced more minimal circuits, their effectiveness was comparable to random circuits of the same size, suggesting poor alignment with ground-truth task structure. Thus, we opted not to include mean-ablated circuits in our main results. Detailed examples and ablation configurations are provided in Appendix J.
@@ 6 Two-Phase SQL Generation: Intent First, Grounding Later
We provide additional visualizations demonstrating similar trends across other benchmarks and model configurations in Appendix L.
@@ 7.1 Ablation Validates Robustness–Granularity Tradeoff
Our ablation results reveal that zero-ablation recovers performance significantly better than random circuits, making it more reliable for identifying robust, task-relevant components. In contrast, mean-ablation yields smaller circuits but recovers performance similar to random baselines, suggesting weaker causal validity. Thus, zero-ablation is better for robustness, while mean-ablation may be useful for identifying compact, interpretable circuits when some loss in reliability is acceptable. Notably, across prompts within all classes, activation patching results vary significantly, even among prompts in the same class, highlighting the inherent variability in the circuits identified (see Appendix F for examples).
@@ Appendix C Dataset details
C.1 Dataset Statistics for CSx_Syn
The full dataset contains 300,000 queries across all variants split into train, validation, and test sets, with 100,000 queries per command set. The average query length using the meta-llama/Llama-3.2-1B-Instruct tokenizer ranges from 12.91 tokens for CS1, 26.52 tokens for CS2, and 41.76 tokens for CS3.
C.2 Dataset Generation Methodology: CSx, CSx_Syn
@@
C.3 Dataset Generation Methodology: CSx_Nat
@@ Appendix E Data set entry characteristics
Dataset Rarity Lexical Density Readab- ility
CS1_Synonyms 0.602 0.580 45.533
CS2_Synonyms 0.525 0.565 29.378
CS3_Synonyms 0.449 0.608 29.850
CS4_Synonyms 0.413 0.579 28.722
CS5_Synonyms 0.443 0.589 25.742
Table 9: Linguistic characteristics of CS*_Synonyms datasets.
@@ Appendix I SAE Circuits Results
Model Dataset Important Heads (Layer1) Recovered Accuracy Component %
@@
Appendix L Additional LogitLens Visualizations

### v4

@@ Abstract
Abir Harrasse*1,5  Philip Quirke*1thanks: * Equal contributions. Correspondence to: philip@withmartian.com.  Clement Neo*2 Dhruv Nathawani3Luke Marks1Amir Abdullah4,1 1Martian  2Apart Research  3NVIDIA  4Thoughtworks  5Mohammed VI Polytechnic University
@@ 1 Introduction
Existing text-to-SQL datasets like Spider (Yu et al., 2019b) and WikiSQL (Zhong et al., 2017) are too complex and noisy for rigorous interpretability analysis. To address this, we introduce TinySQL, a curated dataset that enables controlled analysis of how transformers learn and generate SQL queries. TinySQL progresses through text-to-SQL tasks of increasing complexity, isolating key aspects of the generation process while maintaining structural consistency. Alongside the dataset, we train and release models of 33M, 0.5B, and 1B parameters, providing a comprehensive testbed for interpretability research.
@@
 [^1] Code available at: https://github.com/withmartian/TinySQL,
Datasets, Models and SAEs at: https://huggingface.co/collections/withmartian/tinysql-6760e92748b63fa56a6ffc9f
@@ 2.1 Mechanistic Interpretability Approaches
Beyond mechanistic interpretability, recent work has leveraged synthetic languages and controlled data to probe model reasoning. Chiang and Lee (2022) demonstrate that pre-training on artificial language sequences with explicit and implicit token dependencies improves downstream performance. Allen-Zhu and Li (2023) use CFG-based synthetic grammars to study recursive reasoning, revealing interpretable internal structures and algorithmic attention patterns in generative models. These results highlight the value of structured tasks, and support our use of Text-to-SQL as a semantically grounded, interpretable setting.
@@
Recent work with SAEs (Huben et al., 2023; Templeton et al., 2024) has enabled large-scale analysis of model behavior. Marks et al. (2024) use SAE features to uncover interpretable circuits for tasks like subject-verb agreement, demonstrating the promise of this approach. However, such studies often focus on a single model and task, leaving questions about generalization unresolved. To address this, we introduce multiple Text-to-SQL variants across models, enabling systematic investigation of behavioral variation.
@@ 2.3 Text-to-SQL Datasets
The field evolved from simpler beginnings with WikiSQL (Zhong et al., 2017), which focused on single-table queries from Wikipedia to enable large-scale evaluation. Spider (Yu et al., 2019b) established a foundation for text-to-SQL research by focusing on cross-domain generalization through multiple tables and schemas. Its variants explore different aspects of complexity: Spider-Syn (Gan et al., 2021) tests semantic understanding by replacing schema terms with synonyms, while Spider-Real (Deng et al., 2020) aligns more closely with natural user queries by omitting explicit column names. These datasets collectively enable research into model robustness across synthetic and real-world scenarios.
Recent text-to-SQL datasets have advanced the field by incorporating conversational structure and domain-specific challenges. CoSQL (Yu et al., 2019a) introduces multi-turn interactions grounded in real user dialogues. SParC (Yu et al., 2019c) similarly emphasizes context-dependent SQL generation, where models must resolve references across sequential questions in cross-domain settings. In the medical domain, MIMICSQL (Wang et al., 2020) features naturally varied questions about electronic health records, including paraphrasing and under-specification.
Alongside these realistic benchmarks, recent efforts like Gretel’s Synthetic-Text-To-SQL (Meyer et al., 2024) and SQL-create-context (b mc2, 2023) seek to synthesize datasets that capture naturalistic query patterns while retaining control over generation. These works reflect growing emphasis on bridging the gap between controlled data creation and real-world applicability.
However, these existing datasets prioritize training high-performance models through diverse, complex queries rather than supporting controlled experiments. Even synthetic resources like Sythnetic-Text-To-SQL lack the systematic progression of complexity needed for mechanistic interpretability research. This gap motivates our development of TinySQL, which draws inspiration from machine learning’s long history of using synthetic data to control task complexity and enable clear evaluation. By providing carefully controlled progression of query complexity while maintaining consistent structure, TinySQL bridges the divide between performance-focused datasets and the controlled environment needed for interpretability research.
@@ 3.1 Dataset Structure
Dataset statistics. The complete dataset consists of nine task variants (CS1/2/3, CS1/2/3/4/5_Syn, and CS1_Nat), each containing 100,000 examples generated through a systematic data creation pipeline that ensures consistency with our design principles. The average query length using the meta-llama/Llama-3.2-1B-Instruct tokenizer ranges from 16.48 tokens for CS1_Syn, 38.76 tokens for CS2_Syn, 42.39 tokens for CS3_Syn, 53.73 tokens for CS4_Syn and 61.61 tokens for CS5_Syn, with similar figures for the non synonym variants.
Employing the approach of Mitsopoulou and Koutrika (2025), we compare the English portion of TinySQL to prior Text-to-SQL datasets and find it to be both lexically and semantically more challenging than prior datasets. This is reflected in three key metrics: (a) rarity, the ratio of rare words to content words in a natural language (NL) question Read (2000); (b) lexical density, the ratio of content words to total words Johansson (2008); and (c) Flesch reading ease, a readability measure where lower scores indicate more complex sentences with longer structures and more syllables per word on average Flesch and Gould (1949). As shown in Table 1, TinySQL scores higher on rarity and lexical density, and lower on readability, compared to most previous Text-to-SQL datasets.
Dataset Rarity Lexical Density Readability
CS1_Syn 0.60 0.58 45.5
CS2_Syn 0.52 0.56 29.4
CS3_Syn 0.45 0.61 29.9
CS4_Syn 0.41 0.58 28.7
CS5_Syn 0.44 0.59 25.7
IMDb 0.23 0.43 88.0
Yelp 0.28 0.51 85.0
Academic 0.24 0.54 65.0
Spider 0.21 0.55 89.0
Geo 0.29 0.58 95.0
Restos 0.26 0.56 90.0
MIMICSQL 0.25 0.61 68.0
K-DBQA 0.26 0.59 75.0
BIRD 0.30 0.57 73.0
Atis 0.35 0.46 78.0
Scholar 0.31 0.52 77.0
Advising 0.25 0.44 80.0
Table 1: Dataset Statistics Comparison to Previous Text-to-SQL Datasets
For details on our dataset generation methodology, including schema design, query construction, and instruction templating for CSx, CSx_Syn, and CSx_Nat, please refer to Appendices C.1 and C.2.
@@ 4.4 Sparse Circuit Identification: Method and Results
[…]
@@
To isolate task-relevant features, we follow the four-step procedure in Algorithm 1. We evaluate each circuit’s fidelity using zero-ablation and report recovered accuracy in Table 3, with additional results on other combinations in Appendix. I.
@@
Table 3: Recovered accuracy and layer-wise head selections for circuits identified via high-AUC SAE features. Each layer column is split into two rows per model (first row for attention heads and second for MLP neurons). Full results are in Appendix. H.
@@ 5.1 Identifiability via SAE-EAP Overlap
Model Dataset
Layer
0
Layer
1
@@ 5.2 Local Minimality of Circuits
Model Dataset
Percentage of
Model Used
@@
We also explored an alternative circuit extraction strategy based on mean ablation . While this approach produced more minimal circuits, their effectiveness was comparable to random circuits of the same size, suggesting poor alignment with ground-truth task structure. Thus, we opted not to include mean-ablated circuits in our main results. Detailed examples and ablation configurations are provided in Appendix. J.
@@ 6 Two-Phase SQL Generation: Intent First, Grounding Later
We provide additional visualizations demonstrating similar trends across other benchmarks and model configurations in Appendix. M and a detailed logit lens analysis illustrating the two-phase token emergence across layers in Appendix. L
@@ 7.1 Ablation Validates Robustness–Granularity Tradeoff
Our ablation results reveal that zero-ablation recovers performance significantly better than random circuits, making it more reliable for identifying robust, task-relevant components. In contrast, mean-ablation yields smaller circuits but recovers performance similar to random baselines, suggesting weaker causal validity. Thus, zero-ablation is better for robustness, while mean-ablation may be useful for identifying compact, interpretable circuits when some loss in reliability is acceptable. Notably, across prompts within all classes, activation patching results vary significantly, even among prompts in the same class, highlighting the inherent variability in the circuits identified (see Appendix.F for examples).
@@ Appendix C Dataset details
C.1 Dataset Generation Methodology: CSx, CSx_Syn
@@
C.2 Dataset Generation Methodology: CSx_Nat
@@ Appendix I SAE Circuits Results
Model Dataset Important Heads (Layer0) Recovered Accuracy Component %
@@
Appendix L Logit Lens Analysis Case Study
We performed a comprehensive logit lens analysis on three representative examples from CS2 and CS3 using the BM2 model (Qwen 0.5B, 24 layers) to better understand how SQL generation emerges across transformer layers.
Progressive Token Emergence Pattern: SQL keywords (e.g., SELECT, FROM, WHERE) achieve high confidence (top ranks) in earlier layers (17–21). In contrast, table and column names (e.g., orders, order_id, links) only emerge with high confidence in the final layer (24). Example: In Table 19, across all cases, FROM tokens reach rank 1–2 by layers 21–22, while table names like orders and links remain low until jumping to rank 1 at layer 24.
Logit Trajectory Analysis: A systematic reversal in relative logit strengths is observed. Early layers (0–16): SQL keywords maintain higher logits than table/column names. Later layers: Table/column names surpass SQL keywords in logit values. Example: In Example 1, SELECT has logits around 0.4–0.5 through layers 8–9, while orders starts near 0.05 and only reaches comparable logits (> 0.3) in layer 24.
Token-Type Stratification: Three distinct emergence phases were found:
  • Basic SQL tokens (SELECT, FROM) stabilize by layers 17–21
  • Functional tokens (COUNT, ORDER BY) emerge in mid-layers
  • Context-specific identifiers (table/column names) resolve in the final layer
This supports a hierarchical decoding pattern: from general syntax to specific schema grounding.
Cross-Example Consistency: The same patterns hold across diverse SQL operations: COUNT aggregation (Example 1), MIN functions (Example 2), and ORDER BY clauses (Example 3). Schema-specific tokens always emerge last (layer 24), while SQL structure stabilizes earlier.
Examples Analyzed
  • Example 1
    Instruction: In orders, list total count amount and instances of duration
    Context: CREATE TABLE orders (duration TIME, amount INT)
    Response: SELECT COUNT(amount) AS COUNT_amount, COUNT(duration) AS COUNT_duration FROM orders
  • Example 2
    Instruction: Search for least duration in links beginning with the most duration
    Context: CREATE TABLE links (price INT, duration TIME)
    Response: SELECT MIN(duration) AS MIN_duration FROM links
  • Example 3
    Instruction: Show me the duration and size from the orders table ordered by duration in ascending order
    Context: CREATE TABLE orders (duration TIME, size INT)
    Response: SELECT duration, size FROM orders ORDER BY duration ASC
The results are summarized in Table 19
Summary Table: Layer Where Token First Achieves Top Rank
Token Type Example 1 Example 2 Example 3
FROM Layer 21 Layer 21 Layer 22
Aggregation (MIN, COUNT) Layer 17 Layer 21 N/A
AS Layer 24 Layer 24 N/A
GROUP BY N/A N/A Layer 17
Table Name (links, orders) Layer 24 Layer 24 Layer 24
Column Name Layer 24 Layer 24 Layer 24
Table 19: Layer where each token type first achieves top rank.
Logit Progression Pattern
  • Early Layers (0–16): SQL keywords dominate with a positive logit advantage
  • Mid Layers (17–21): SQL structure tokens reach high ranks, schema tokens remain low
  • Final Layer (24): Schema-specific tokens (e.g., table/column names) reach top logit values
This progression supports the hypothesis that transformer layers follow a hierarchical reasoning process, from abstract SQL syntax to grounded schema resolution.
We also present in Table 20, a detailed analysis of the SQL Keyword COUNT ranks across layers compared to the table name orders ,
Layer COUNT Rank orders Rank
1 392 43497
2 832 71069
3 1032 67813
4 5652 83330
5 7892 118378
6 4532 85341
7 2209 112749
8 3505 136323
9 2101 144139
10 2065 136692
11 4047 136514
12 2676 132739
13 384 127971
14 92 108117
15 124 76418
16 1 79962
17 1 33234
18 1 37628
19 1 25178
20 1 24058
21 1 7664
22 2 2430
23 2 166
24 1 1
Table 20: Token rank comparison between COUNT and orders across layers for Example 1.
Appendix M Additional LogitLens Visualizations